Pybus, Inc. is considering issuing bonds that will mature in 15 years with an annual coupon rate of 8 percent. Their par value will be $1,000​, and the interest will be paid semiannually. Pybus is hoping to get a AA rating on its bonds​ and, if it​ does, the yield to maturity on similar AA bonds is 12 percent. ​ However, Pybus is not sure whether the new bonds will receive a AA rating. If they receive an A​ rating, the yield to maturity on similar A bonds is 13 percent. What will be the price of these bonds if they receive either an A or a AA​ rating? a. The price of the Pybus bonds if they receive a AA rating will be ​$_____?
Answer #1

Price of bond today is equal to the present value of all future coupon payments and the principal amount

At AA rating,

= 1,000*8%*1/2*PVAF(6%, 30 periods) + 1,000*PVF(6%, 30 periods)

= 40*13.7648 +1,000*0.1741

= $724.69

At A rating,

Price = 1,000*8%*1/2*PVAF(6.5%, 30 periods) + 1,000*PVF(6.5%, 30 periods)

= 40*13.0587+1,000*0.1512

= $673.55